OpenStudy (anonymous):

1,000,000=500,000/(1+i)^2 solve for i

OpenStudy (andriod09):

Heyo Connor. :D Tell me, what's the first thing you should do with this equation?

OpenStudy (andriod09):

have you heard of PEMDAS before?

OpenStudy (andriod09):

Then tell me what PEMDAS means, I'll be here if you can't remember all of it. :)

OpenStudy (anonymous):

parenthesis, exponent, multiply, divide, add, subtract

OpenStudy (andriod09):

That's right, so then using pemdas, what do you do first in your equation of: \[1,000,000 = 500,000\div(1+i)^2\]

OpenStudy (anonymous):

get rid of the exponent

OpenStudy (andriod09):

Tell me, how would you do that?

OpenStudy (anonymous):

raise both sides by 1/2

OpenStudy (anonymous):

hello?

OpenStudy (campbell_st):

ok... so start by swapping (1 + i)^2 and 1000000 so \[(1 + i)^2 = \frac{500000}{1000000} = \frac{1}{2}\] so if \[(1 + i)^2 = \frac{1}{2}\] take the square root of both sides then subtract 1 to get the value of 1

OpenStudy (campbell_st):

oops value of i